A sample of 100 Grade 9 students’ ages have been obtained to estimate the mean age of all Grade 9 students. ¯x=16 years and the population variance is 0.75 years. Find the 95% confidence interval for µ.

1
Expert's answer
2022-04-12T17:30:33-0400

"95\%CI=(16-1.96\frac{0.75}{\sqrt{100}}, 16+1.96\frac{0.75}{\sqrt{100}})=(15.85, 16.15)."